How many milliliters of 3 F KOH can be prepared from 100 grams of KOH?


1
Expert's answer
2021-02-28T06:19:16-0500

CM = n/V

V = n / CM

n = m/M

M (KOH) = 56.11 g/mol

n (KOH) = 100 / 56.11 = 11.78 mol

V (KOH) = 11.78 / 3 = 3.93 L = 3 930 ml